OpenStudy (anonymous):

12) Brad wanted to find three consecutive whole numbers that add up to 81. He wrote the equation (n - 1) + n + (n + 1) = 81What does the n stand for?

OpenStudy (anonymous):

the second of the three consecutive numbers.

OpenStudy (anonymous):

the numbers are 26,27,28

OpenStudy (anonymous):

n = 27

OpenStudy (gw2011):

Let n=the first number Let n+1=the second number Let n+2=the third number n+(n+1)+(n+2)=81 n+n+1+n+2=81 3n+3=81 3n=78 n=26 This is the first number n+1=26+1=27 This is the second number n+2=26+2=28 This is the third number Therefore, the three consecutive numbers are 26, 27, 28